# Unifying Multimodal Large Language Model Capabilities and Modalities via Model Merging

### v1

@@ 4 Methodology
WUDI merging defines a loss between the merged vector $\boldsymbol{\tau}_{m}$ and task vectors $\boldsymbol{\tau}_{i}$ to learn the merged vector. Building on this foundation, we propose improved task vector optimization.
For full fine-tuning.
@@
Task vectors contain extreme redundancy and noise that create mutual interference, impeding effective merge vector optimization. To address this issue, we propose reducing inter-task interference through low-rank approximation. First, we calculate the average task vector $\bar{\boldsymbol{\tau}}_{l}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{\tau}_{i,l}$ and use it to center task vectors [15]. Next, we perform SVD to isolate core task-specific knowledge from noise present in the top and lower singular vectors.
$\mathrm{SVD}(\boldsymbol{\tau}_{i,l}-\bar{\boldsymbol{\tau}}_{l})=U\Sigma V^{\top},\text{ where }U\in\mathbb{R}^{m\times r},\Sigma\in\mathbb{R}^{r\times r},V\in\mathbb{R}^{n\times r}.$ (3)
@@
We then apply low-rank approximations to eliminate redundant noise, where $U_{1:k},\Sigma_{1:k},V_{1:k}^{\top}$ represent the top-$k$ singular components. Moreover, we find that substituting $\Sigma_{1:k}V_{1:k}^{\top}$ for the task vector $\boldsymbol{\tau}_{i,l}$ as the input subspace $\boldsymbol{x}_{i,l}$ allows us to discard secondary row space information, focusing only on the column feature space. Thus, we can optimize $\boldsymbol{\tau}_{m,l}$ via gradient decent on the loss:
$\min_{\boldsymbol{\tau}_{m,l}}\mathcal{L}_{l}=\sum_{i=1}^{n}\frac{1}{\left\|\boldsymbol{\tau}_{i,l}\right\|_{F}^{2}}\left\|(\boldsymbol{\tau}_{m,l}-U_{1:k}\Sigma_{1:k}V_{1:k}^{\top}-\bar{\boldsymbol{\tau}}_{l})(\Sigma_{1:k}V_{1:k}^{\top})^{\top}\right\|_{F}^{2}.$ (4)
@@
By truncating singular values, we preserve critical features $V_{1:k}^{\top}$, which is similar to selecting principal components in PCA [1]. This results in more accurate input $\boldsymbol{x}_{i,l}$ estimations.
For LoRA fine-tuning.
@@
Figure 3: $\boldsymbol{\tau}_{m}$ tends to take shortcuts by increasing its magnitude to encourage orthogonality.
@@
Figure 4: Frobenius norm of the merged vector changes during optimization (average by layers).
@@
The inherent low-rank nature of LoRA fine-tuning presents unique optimization challenges for the merge vector. When optimizing $\boldsymbol{\tau}_{m,l}$, gradients become effective only in directions corresponding to non-zero singular values of $\boldsymbol{\tau}_{i,l}$, while approaching zero in other directions (null space). This constraint limits parameter update freedom, preventing $\boldsymbol{\tau}_{m,l}$ from properly exploring the parameter space. We observe that $\boldsymbol{\tau}_{m,l}$ tends to take shortcuts by increasing its magnitude to minimize loss. This occurs because the merge vector must simultaneously accommodate multiple task vectors in different directions. Without constraints, Eq. 1 encourages orthogonality by increasing the length of the merge vector (see Fig. 4). When added to the base model, such large-norm task vectors cause deviation from the original distribution, resulting in collapsed language ability [90]. To address this, (i) we adopt SGD instead of Adam, as it better escapes local optima in flat regions and offers greater stability with sparse gradients. (ii) We propose a direct low-rank approximation of $\boldsymbol{\tau}_{i,l}$ without centering, which helps reduce the Frobenius norm of task vectors. (iii) We also introduce initializing the merged vector with the mean of task vectors to mitigate the issue of excessive merge vector magnitude. As shown in Fig. 4, our approach maintains a relatively consistent norm throughout the optimization process while minimizing loss.
@@ 5.1 Benchmark Details
Checkpoint Construction
We select two types of vision-language models representing different scales and usage scenarios: InternVL2.5-1B-Instruct [13], a lightweight aligned model, and Qwen2-VL-7B-Base [77], a general-purpose foundation model. “Instruct” denotes the instruction-following version of a pre-trained model, which is the variant most commonly released. In contrast, Qwen2-VL-Base is one of the few publicly available pre-trained model versions. For InternVL2.5-1B-Instruct, we perform full fine-tuning with a learning rate of 4e-5 and a warmup ratio of 3e-2. For Qwen2-VL-7B-Base, we apply LoRA fine-tuning with a rank of 8, a learning rate of 1e-5, and a warmup ratio of 1e-1. Both models are trained for one epoch using a cosine learning rate scheduler. Different training strategies and scales help evaluate the generalizability of merging methods.
For modality merging, we follow prior work [7] and select Vicuna-7B-v1.5 [97] as the shared LLM. The vision-language model uses CLIP-ViT-L-336px [66] as the image encoder, paired with an MLP projection as the connector. The audio-language model adopts BEATs-Iter3+ [9] as the audio encoder, with a Q-Former incorporating 32 query tokens as the connector. The video-language model employs LanguageBind [98] as the video encoder, also using an MLP connector. See App. D for details.
Training Data
Following [13], we collect a broader range of domain-specific data, divided into VQA, Geometry, Chart, OCR, and Grounding tasks. The datasets used are summarized in Table 1. For effective supervised fine-tuning, we gather at least 100k public dataset samples for each task, ensuring maximum diversity wherever possible. We process all data into the instruction tuning format. Specifically, for grounding tasks, we map coordinates to the [0,1000) range and add the special token notation <|box_start|><|box_end|> [77]. During Qwen2-VL-Base fine-tuning, we observe that Chinese datasets consistently degraded performance, possibly due to lower data quality or because evaluation benchmarks are primarily in English. Consequently, we use only English datasets for instruction tuning of Qwen2-VL-Base. InternVL2.5-Instruct, already possessing multilingual instruction-following capabilities, is fine-tuned using all available data.
Table 1: Summary of task-specific datasets, with their corresponding sizes and languages.
@@
VQA 588K GQA (en) [34], VQAv2 (en) [29], OKVQA (en) [57], LLaVA-Instruct (zh) [48], CogVLM-Singleround (en&zh) [78], CogVLM-Multiround (en&zh) [78]
Geometry 190K GeoQA+ (zh) [6], G-LLaVA (en) [23]
Chart 218K ChartQA (en) [58], DVQA (en) [37]
OCR 238K OCRVQA (en) [62], TextCaps (en) [67], SynthDoG (en) [40], LLaVAR (en) [96], ST-VQA (en) [5], TextVQA (en) [68], DocVQA (en) [59], DeepForm (en) [72], KLC (en) [69], TabFact (en) [10]
Grounding 135K RefCOCO (en) [92, 55], VG (en) [41]
Evaluation Benchmark
Current benchmarks [50, 42, 8, 22] predominantly evaluate a model’s overall performance but provide limited insights into specific capabilities. Therefore, we carefully select specialized datasets to evaluate distinct abilities across tasks. (i) For VQA, we utilize VizWiz [30] and GQA [34] to assess general visual question answering proficiency. (ii) For Geometry, we incorporate multiple challenging subsets: “geometry reasoning”, “algebraic reasoning” and “geometry problem solving” from MathVista [51], complemented by “metric geometry - angle”, “metric geometry - area”, “metric geometry - length” and “solid geometry” from MATH-Vision [76]. (iii) For Chart, we employ ChartQA [58], which tests reasoning and interpretation ability with charts and graphs. (iv) For OCR, our evaluation suite includes TextVQA [68] and OCRVQA [62]. (v) For Grounding, we implement referring expression comprehension using RefCOCO [38], RefCOCO+ [38], and RefCOCOg [55], which require models to identify specific objects in images based on natural language descriptions. All evaluation results are obtained using the VLMEvalKit [20] and LMMs-Eval [95] libraries under the same settings to ensure fair comparison. When evaluating MathVista and MATH-Vision benchmarks, we utilize the GPT-4o-mini API to extract answers from the output. All experiments are conducted using 8$\times$ NVIDIA V100 GPUs.
For Omni-language models, we select audio-visual question answering task, which involves answering questions about visual objects, sounds, and their relationships in videos. This task requires multimodal understanding and spatio-temporal reasoning over audio-visual scenes. AVQA [89] targets real-world objects and activities. MUSIC-AVQA [44] specifically focuses on musical performances.
5.2 Merging Details
Following Task Arithmetic [35], we employ a single coefficient $\lambda$ to scale the merged vector before adding it to the base model. For all model merging methods, we determine the optimal merging coefficient $\lambda$ by searching within the range [0.1, 0.3, 0.5, 0.7, 1.0, 1.5]. In our implementation, we use the Adam optimizer with a learning rate of 1e-5 for InternVL while applying the SGD optimizer with a learning rate of 1e-4 for QwenVL. The number of optimization iterations is set to 300. Consistent with prior work [14], we apply our method exclusively to the linear layer in the model.
@@ 5.3 Experimental Results
Weight Average [82] 29.96 54.89 49.60 18.42 71.64 74.54 41.86 52.62 45.29 52.39 49.12
Task Arithmetic [35] 30.67 56.34 45.36 21.05 72.88 76.26 43.39 74.90 68.15 72.75 56.18
TIES Merging [84] 30.63 56.48 44.50 23.68 72.28 76.29 44.01 76.01 68.45 73.65 56.70
TA w/ DARE [93] 30.61 56.48 48.45 21.05 73.08 76.30 43.03 74.94 68.07 73.02 56.50
TIES w/ DARE [93] 30.65 56.11 43.85 27.63 72.72 76.19 43.33 75.10 68.48 73.55 56.76
TSV Merging [24] 31.15 56.67 52.45 28.95 70.56 75.66 45.38 65.19 58.51 59.17 54.37
Iso-C [56] 28.21 55.36 48.96 21.05 70.56 69.34 46.51 72.72 66.56 68.50 54.78
WUDI Merging [14] 31.02 56.96 53.03 17.11 69.19 75.95 46.12 76.06 70.14 74.48 57.00
WUDI v2 (Ours) 30.97 57.13 54.48 21.05 68.72 76.01 46.35 75.97 69.72 73.94 57.44
@@
Weight Average [82] 41.47 57.33 50.21 34.21 59.56 81.09 57.85 80.72 65.37 77.68 60.55
Task Arithmetic [35] 40.52 62.31 40.36 26.31 79.67 81.09 59.50 75.96 61.33 75.85 60.29
TIES Merging [84] 41.38 59.08 46.87 34.21 67.24 81.42 58.53 80.63 65.36 77.65 61.24
TA w/ DARE [93] 40.64 62.38 40.67 26.31 79.76 81.04 59.34 75.83 61.41 75.80 60.32
TIES w/ DARE [93] 41.63 59.96 45.72 35.53 70.68 81.53 59.63 80.73 65.65 77.77 61.88
TSV Merging [24] 41.43 57.31 51.05 34.21 59.44 81.25 57.81 80.71 65.34 77.76 60.63
Iso-C [56] 12.31 13.44 39.96 27.63 2.80 30.05 6.12 53.68 38.96 41.90 26.69
WUDI Merging [14] 37.19 56.45 42.96 27.63 67.84 79.92 65.56 76.25 60.72 71.99 58.65
WUDI v2 (Ours) 41.61 61.16 48.66 40.79 74.08 81.54 60.06 80.92 65.90 78.24 63.30
@@
Individual Modalities Merging Methods Online Composing
Datasets Vision Audio Video
Weight
Average
[82]
Task
Arithmetic
[35]
Ties
Merging
[84]
TSV
Merging
[24]
Iso-C
[56]
WUDI
Merging
[14]
WUDI v2
NaiveMC
[7]
DAMC
[7]
Individual models.
As shown in Table 2 and Table 3, models obtained through SFT demonstrate improvements across their respective tasks. Notably, Qwen2-VL-Base, which initially lacks instruction-following capabilities, learns the correct response format after SFT. For the already comprehensive InternVL2.5-Instruct model, individual models enhance specific capabilities but sacrifice generalization performance [25, 94]. We observe that OCR and Grounding tasks exhibit complementary benefits, indicating task similarity. In the table, the “-” indicates instances where the model failed to output in the correct grounding format, i.e., normalized coordinate boxes.
Capability merging.
Task vectors from individual models effectively capture specialized capabilities, enabling successful model merging. By combining these strengths, model merging outperforms expert MLLMs in their respective tasks. For example, the merged Qwen2-VL significantly outperforms individual models on Geometry task (51.05 and 40.79 vs. 42.50 and 28.95) and Chart task (79.76 vs. 61.08). Similar improvements are also observed for OCR and Grounding tasks. For InternVL2.5-Instruct, we conduct mixture training by combining all task-specific training data for SFT. For Qwen2-VL-Base, we directly use Qwen2-VL-Instruct as the upper bound for mixture training, given its extensive prior SFT with diverse datasets. Notably, our best model merging methods closely match or even surpass mixture training and instruct versions. These results demonstrate that model merging potentially surpasses multi-task learning while providing a scalable solution for creating high-performing MLLMs with reduced computational cost and time.
Categorization of merging methods.
Different model merging methods demonstrate distinct characteristics. Linear interpolation of task vectors, while neglecting parameter conflicts, achieves robust yet only moderate performance. Sparsification-based methods assume that pruning redundant parameters improves outcomes. TIES struggles to control sparsity levels, often underperforming compared to task arithmetic. DARE consistently delivers plug-and-play improvements through rescaling operations that approximate the original embedding. SVD-based methods’ performance correlates with task vectors’ spectral properties. For example, Iso-C fails on Qwen2-VL because the LoRA-tuned task vectors are already low-rank, and averaging singular values further reduces their Frobenius norm, creating instability in LLMs. Even increasing $\lambda$, as recommended in their paper, only marginally improves results. TSV merging excels in modality merging because its orthogonalization mitigates modal conflicts, but delivers ordinary performance in multi-task settings. In contrast, our approach achieves superior average results across various scenarios, benefiting from more accurate task vector estimation and more stable optimization.
Modality merging.
As shown in Table 4, merging methods effectively integrate information from three modalities, outperforming models trained on individual vision, audio, or video inputs. This highlights the complementary nature of modal information and its potential for merging. Online composing dynamically merges activations in the LLM from different modalities during inference, requiring separate parameter storage for each modality (i.e., 3$\times$ static merging). NaiveMC [7] performs simple activation averaging, while DAMC [7] decouples parameters during training to reduce modal interference. Notably, the best merging method even outperforms these online composition methods. Advancing Omni models through model merging offers a promising direction for future research.
Improved task vector optimization.
@@
Our method enhances task vector optimization stability, achieving optimal results. In Table 5, we evaluate each component’s contribution to overall performance. Starting with WUDI merging, we incrementally add one component at a time, reporting performance for both PEFT model merging (Qwen2-VL) and modality merging (Vicuna-7B). Replacing Adam with SGD alone does not necessarily improve performance; however, when combined with initializing the merged vector using the mean of task vectors, we observe a significant 4.43% improvement. Low-rank approximation further enhances performance, demonstrating its effectiveness in preserving critical knowledge from task vectors while maintaining the stability of the Frobenius norm.
Computational requirements.
Table 6: Model merging vs. Data mixing.
Methods Solving Time GPU Memory
InternVL2.5-1B (Ours) 0.22h 2.62GB
InternVL2.5-1B (Mixed) 25.38h 240GB
Qwen2-VL-7B (Ours) 3.78h 21.97GB
Qwen2-VL-7B (Mixed) 24.56h 256GB
As illustrated in Table 6, we compare the solving time and GPU memory usage of our approach against mixture training. Our approach optimizes the merged vector over 300 iterations while incurring minimal computational overhead and requiring significantly less GPU memory than data-based training. This efficiency is achieved through layer-by-layer optimization without requiring training data. Our results confirm that the proposed method is computationally efficient and highly scalable on devices with modern GPUs, facilitating the rapid development of new models based on existing ones.
Actual checkpoints from Hugging Face.
To evaluate the practicality of model merging in communities, we collect fine-tuned models released by different developers on Hugging Face. Our collection includes a model specialized in math reasoning via multimodal reinforcement learning [^1] https://huggingface.co/lmms-lab/Qwen2-VL-7B-GRPO-8k, a personalized model for the Pokemon domain [^2] https://huggingface.co/hiyouga/Qwen2-VL-7B-Pokemon, a model focused on converting PDF documents into text [^3] https://huggingface.co/allenai/olmOCR-7B-0225-preview, and a model with OCR and VQA capabilities in Vietnamese [^4] https://huggingface.co/erax-ai/EraX-VL-7B-V1.0. As shown in Table 7, various merging methods achieve performance that surpasses that of the individual models. These results demonstrate that model merging can effectively integrate knowledge from diverse models to construct a more robust system, underscoring its promising practical applications.
@@
Weight Average [82] 43.74 61.30 43.96 26.63 76.40 82.11 71.61 89.54 82.18 86.43 66.39
Task Arithmetic [35] 41.57 60.95 42.99 23.68 75.28 81.95 71.78 87.72 81.60 85.63 65.32
TIES Merging [84] 44.17 60.54 42.52 27.95 75.48 82.40 71.09 90.06 83.52 86.44 66.42
TA w/ DARE [93] 43.33 61.15 44.37 26.95 76.48 82.93 72.00 88.93 82.79 86.07 66.50
TIES w/ DARE [93] 44.37 60.78 44.37 27.63 76.04 82.61 70.93 89.40 82.93 86.77 66.58
TSV Merging [24] 43.73 61.40 43.54 27.94 76.44 83.15 71.65 88.53 82.25 86.41 66.50
Iso-C [56] 43.99 61.34 40.91 22.37 76.96 83.33 71.55 87.74 82.10 85.27 65.56
WUDI Merging [14] 41.39 60.11 44.20 21.05 74.36 80.78 71.12 87.96 81.50 85.48 64.80
WUDI v2 (Ours) 43.76 61.29 44.68 27.63 76.24 82.97 71.48 89.56 82.97 86.42 66.70
@@
6 Limitation and Future Work
Due to resource constraints, our experiments were limited to models of 7B parameters. The public datasets we collected may contain lower-quality data that hasn’t undergone rigorous cleaning. Future work will explore multilingual or reasoning-focused MLLM merging, incorporating visual chain-of-thought datasets [91, 46] to support expert reasoning models. For evaluation, we plan to develop new benchmarks specifically designed to assess the reasoning capabilities of MLLMs.

### v2

@@ 4 Methodology
Eq.˜1 defines a loss between the merged vector and the task vectors. However, data-free optimization often suffers from instability and convergence issues. To address this, we propose OptMerge, a novel method that improves task vector optimization. Specifically, our approach accommodates both full fine-tuning and LoRA fine-tuning scenarios, as they yield model parameters with distinct properties (e.g., low-rank sparsity and varying optimization difficulty). These differences naturally necessitate tailored merging strategies, as detailed in Sec.˜4.1 and Sec.˜4.2.
4.1 Merging Full Fine-tuning Models
@@
Task vectors contain significant redundancy and noise, leading to mutual interference during merging. Redundancy stems from different tasks re-learning shared foundational skills, while noise reflects non-essential parameter updates. Directly adding task vectors amplifies these issues, hindering effective merge vector optimization. To address this issue, we propose reducing inter-task interference through low-rank approximation. First, we calculate the average task vector $\bar{\bm{\tau}}_{l}=\frac{1}{n}\sum_{i=1}^{n}\bm{\tau}_{i,l}$ and use it to center task vectors (Choi et al., 2024). Next, we perform SVD to isolate core task-specific knowledge from noise present in the top and lower singular vectors [^1] We optimize only the model’s linear layers, representing each task vector $\bm{\tau}_{i,l}$ as an $m\times n$ matrix. The remaining layers are merged by simple parameter averaging..
$\mathrm{SVD}(\bm{\tau}_{i,l}-\bar{\bm{\tau}}_{l})=U\Sigma V^{\top},\text{ where }U\in\mathbb{R}^{m\times r},\Sigma\in\mathbb{R}^{r\times r},V\in\mathbb{R}^{n\times r}.$ (3)
@@
We then apply low-rank approximations to eliminate redundant noise, where $U_{1:k},\Sigma_{1:k},V_{1:k}^{\top}$ represent the top-$k$ singular components. Moreover, we find that substituting $\Sigma_{1:k}V_{1:k}^{\top}$ for the task vector $\bm{\tau}_{i,l}$ as the input subspace $\bm{x}_{i,l}$ allows us to discard secondary row space information, focusing only on the column feature space. Thus, we can optimize $\bm{\tau}_{m,l}$ via gradient descent on the loss:
$\min_{\bm{\tau}_{m,l}}\mathcal{L}_{l}=\sum_{i=1}^{n}\frac{1}{\left\|\bm{\tau}_{i,l}\right\|_{F}^{2}}\left\|(\bm{\tau}_{m,l}-U_{1:k}\Sigma_{1:k}V_{1:k}^{\top}-\bar{\bm{\tau}}_{l})(\Sigma_{1:k}V_{1:k}^{\top})^{\top}\right\|_{F}^{2}.$ (4)
@@
By truncating singular values, we preserve critical features $V_{1:k}^{\top}$, which is similar to selecting principal components in Principal Components Analysis (PCA) (Abdi & Williams, 2010). This yields more accurate estimates of $\bm{x}{i,l}$ than using $(\bm{\tau}_{i,l})^{\top}$.
@@
Figure 3: When optimizing Eq.˜1, $\bm{\tau}_{m}$ tends to take shortcuts by increasing its magnitude to achieve orthogonality.
@@
Figure 4: We plot the progression of the Frobenius norm of the merged vector during optimization (average by layers).
4.2 Merging LoRA Fine-tuned Models
@@
The inherent low-rank nature of LoRA fine-tuning presents unique optimization challenges for the merge vector. When optimizing $\bm{\tau}_{m,l}$, gradients become effective only in directions corresponding to non-zero singular values of $\bm{\tau}_{i,l}$, while approaching zero in other directions (null space). This constraint limits parameter update freedom, preventing $\bm{\tau}_{m,l}$ from properly exploring the parameter space. We observe that $\bm{\tau}_{m,l}$ tends to take shortcuts by increasing its magnitude to minimize loss. This occurs because the merge vector must simultaneously accommodate multiple task vectors in different directions. Without constraints, Eq.˜1 achieves orthogonality by increasing the length of the merge vector (see Fig.˜4). When added to the base model, such large-norm task vectors cause deviation from the original distribution, resulting in collapsed language ability.
To address these challenges, we introduce a set of practical techniques: (1) We replace Adam with SGD, which better escapes flat local optima and offers greater stability under sparse gradients. Notably, SGD provides implicit regularization (Smith et al., 2021; Wang et al., 2022), constraining task vector optimization and navigating flat regions induced by null spaces. (2) We apply a direct low-rank approximation to $\bm{\tau}_{i,l}$ using truncated $\mathrm{SVD}(\bm{\tau}_{i,l})\approx U_{1:k}\Sigma_{1:k}V_{1:k}^{\top}$ without centering. The Frobenius norm equals the sum of squared singular values $\|\bm{\tau}_{i,l}\|_{F}^{2}=\sum_{j=1}^{r}\sigma_{j}^{2}$. After truncation, we drop the tail energy $\sum_{j>k}\sigma_{j}^{2}$ and thus reduce the norm. (3) We also introduce initializing the merged vector with the mean of task vectors to mitigate the issue of excessive merge vector magnitude. As shown in Fig.˜4, our approach maintains a relatively consistent norm throughout the optimization process while minimizing loss successfully.
@@ 5 MLLMs Merging Benchmark
We begin by detailing the benchmark, including its checkpoints, datasets, and evaluation protocols, as well as the implementation of merging algorithms (Sec.˜5.1). Next, Sec.˜5.2 presents extensive experiments that empirically validate the benchmark, and summarizes the key findings.
@@ 5.1 Benchmark Details
Checkpoint construction. To cover two practical scenarios, namely fine-tuning base models and fine-tuning instruction-tuned models, we select two models that differ in intended use: InternVL2.5-1B-Instruct (Chen et al., 2024c), a lightweight model aligned for instruction following, and Qwen2-VL-7B-Base (Wang et al., 2024b), a general foundation model. Qwen2-VL-7B-Base is among the few publicly available pretrained models, so we use it for the base-model scenario. These choices span different training strategies and scales, enabling a broad assessment of merging methods.
For modality merging, we select Vicuna-7B-v1.5 (Zheng et al., 2023) as the shared LLM. The vision-language model uses CLIP-ViT-L-336px (Radford et al., 2021) as the image encoder, paired with an MLP projection as the connector. The audio-language model adopts BEATs-Iter3+ (Chen et al., 2023) as the audio encoder, with a Q-Former as the connector. The video-language model employs LanguageBind (Zhu et al., 2023) as the video encoder. See App. C for details.
Training data. We collect a broader range of domain-specific data, divided into VQA, Geometry, Chart, OCR, and Grounding tasks. The datasets used are summarized in Table˜1. For effective supervised fine-tuning, we gather at least 100k public dataset samples for each task, ensuring maximum diversity wherever possible. We process all data into the instruction tuning format.
Table 1: Summary of collected training datasets, with their corresponding sizes and languages.
@@
VQA 588K GQA (en) (Hudson & Manning, 2019), VQAv2 (en) (Goyal et al., 2017), OKVQA (en) (Marino et al., 2019), LLaVA-Instruct (zh) (Liu et al., 2024a), CogVLM-Singleround (en&zh) (Wang et al., 2024c), CogVLM-Multiround (en&zh) (Wang et al., 2024c)
Geometry 190K GeoQA+ (zh) (Cao & Xiao, 2022), G-LLaVA (en) (Gao et al., 2023)
Chart 218K ChartQA (en) (Masry et al., 2022), DVQA (en) (Kafle et al., 2018)
OCR 238K OCRVQA (en) (Mishra et al., 2019), TextCaps (en) (Sidorov et al., 2020), SynthDoG (en) (Kim et al., 2022), LLaVAR (en) (Zhang et al., 2023), ST-VQA (en) (Biten et al., 2019), TextVQA (en) (Singh et al., 2019), DocVQA (en) (Mathew et al., 2021), DeepForm (en) (Svetlichnaya, 2020), KLC (en) (Stanisławek et al., 2021), TabFact (en) (Chen et al., 2019)
Grounding 135K RefCOCO (en) (Yu et al., 2016; Mao et al., 2016), VG (en) (Krishna et al., 2017)
[…]
For Omni-language models, we assess Audio-VQA, which requires multimodal understanding and spatiotemporal reasoning about visual objects, sounds, and their relationships in videos.
Merging details. Following Task Arithmetic (Ilharco et al., 2023), we employ a single coefficient $\lambda$ to scale the merged vector before adding it to the base model. For all model merging methods, we determine the optimal merging coefficient $\lambda$ by searching within the range [0.1, 0.3, 0.5, 0.7, 1.0, 1.5]. In our implementation, the rank size $k$ in Eq.˜4 is simply defined as the rank of each task vector divided by the number of tasks (i.e., 5). We use the Adam optimizer with a learning rate of 1e-5 for InternVL while applying the SGD optimizer with a learning rate of 1e-4 for QwenVL. The number of optimization iterations is set to 300. We apply our method exclusively to the linear layer in the model.
@@ 5.2 Experimental Results
Capability merging. As shown in Tables˜2 and 3, merging individually specialized models outperforms expert MLLMs on their target tasks. For example, the merged Qwen2-VL achieves 51.05 and 40.79 on Geometry (vs. 42.50 and 28.95 for individual models) and 79.76 on Chart (vs. 61.08). We observe similar gains for OCR and Grounding, with complementary benefits between these tasks. For InternVL2.5-Instruct, we conduct mixture training by combining all task-specific training data for SFT. For Qwen2-VL-Base, we directly use Qwen2-VL-Instruct as the upper bound for mixture training, given its extensive prior SFT with diverse datasets. Notably, our best model merging methods closely match or even surpass mixture training and instruct versions. These results demonstrate that model merging potentially surpasses multi-task learning, while providing a scalable solution for creating high-performing MLLMs with reduced computational cost.
Categorization of merging methods. Different merging methods exhibit distinct behaviors. Linear interpolation of task vectors, while ignoring parameter conflicts, is robust but only moderately effective. Sparsification-based methods such as TIES struggle to control sparsity and often underperform relative to task arithmetic. DARE reliably provides plug-and-play gains through simple rescaling. SVD-based methods are sensitive to the spectral structure of task vectors. For example, Iso-C fails on Qwen2-VL because the LoRA-tuned task vectors are already low-rank, and averaging singular values further reduces their Frobenius norm, creating instability in LLMs. Even increasing $\lambda$, as recommended in their paper, only marginally improves results. TSV merging excels in modality merging because its orthogonalization mitigates modal conflicts, but delivers ordinary performance in multi-task settings. In contrast, our approach achieves superior average results across various scenarios, benefiting from stable task vector optimization.
@@
Weight Average 29.96 54.89 49.60 18.42 71.64 74.54 41.86 52.62 45.29 52.39 49.12
Task Arithmetic 30.67 56.34 45.36 21.05 72.88 76.26 43.39 74.90 68.15 72.75 56.18
TIES Merging 30.63 56.48 44.50 23.68 72.28 76.29 44.01 76.01 68.45 73.65 56.70
TA w/ DARE 30.61 56.48 48.45 21.05 73.08 76.30 43.03 74.94 68.07 73.02 56.50
TIES w/ DARE 30.65 56.11 43.85 27.63 72.72 76.19 43.33 75.10 68.48 73.55 56.76
TSV Merging 31.15 56.67 52.45 28.95 70.56 75.66 45.38 65.19 58.51 59.17 54.37
Iso-C 28.21 55.36 48.96 21.05 70.56 69.34 46.51 72.72 66.56 68.50 54.78
WUDI Merging 31.02 56.96 53.03 17.11 69.19 75.95 46.12 76.06 70.14 74.48 57.00
OptMerge (Ours) 30.97 57.13 54.48 21.05 68.72 76.01 46.35 75.97 69.72 73.94 57.44
@@
Weight Average 41.47 57.33 50.21 34.21 59.56 81.09 57.85 80.72 65.37 77.68 60.55
Task Arithmetic 40.52 62.31 40.36 26.31 79.67 81.09 59.50 75.96 61.33 75.85 60.29
TIES Merging 41.38 59.08 46.87 34.21 67.24 81.42 58.53 80.63 65.36 77.65 61.24
TA w/ DARE 40.64 62.38 40.67 26.31 79.76 81.04 59.34 75.83 61.41 75.80 60.32
TIES w/ DARE 41.63 59.96 45.72 35.53 70.68 81.53 59.63 80.73 65.65 77.77 61.88
TSV Merging 41.43 57.31 51.05 34.21 59.44 81.25 57.81 80.71 65.34 77.76 60.63
Iso-C 12.31 13.44 39.96 27.63 2.80 30.05 6.12 53.68 38.96 41.90 26.69
WUDI Merging 37.19 56.45 42.96 27.63 67.84 79.92 65.56 76.25 60.72 71.99 58.65
OptMerge (Ours) 41.61 61.16 48.66 40.79 74.08 81.54 60.06 80.92 65.90 78.24 63.30
@@
Improved task vector optimization. Our method enhances task vector optimization stability, achieving optimal results. In Table˜4, we evaluate each component’s contribution to overall performance. Starting with WUDI Merging, we incrementally add one component at a time, reporting performance for both LoRA model merging (Qwen2-VL) and modality merging (Vicuna-7B). Replacing Adam with SGD alone does not necessarily improve performance; however, when combined with initializing the merged vector using the mean of task vectors, we observe a significant 4.43% improvement. Low-rank approximation further enhances performance, demonstrating its effectiveness in preserving critical knowledge from task vectors while maintaining the stability of the Frobenius norm. For full fine-tuned models, Tables˜2 and 6 show average improvements of 0.44% and 1.9% for OptMerge over WUDI Merging, respectively. This highlights the necessity of Eq.˜4 over Eq.˜1 for denoising task vectors and achieving robust merged-vector optimization.
[…]
Individual Modalities Merging Methods Online Composing
Datasets Vision Audio Video
Weight
Average
Task
Arithmetic
Ties
Merging
TSV
Merging
Iso-C
WUDI
Merging
OptMerge
NaiveMC
DAMC
@@
Task Arithmetic 41.57 60.95 42.99 23.68 75.28 81.95 71.78 87.72 81.60 85.63 65.32
TIES Merging 44.17 60.54 42.52 27.95 75.48 82.40 71.09 90.06 83.52 86.44 66.42
TA w/ DARE 43.33 61.15 44.37 26.95 76.48 82.93 72.00 88.93 82.79 86.07 66.50
TIES w/ DARE 44.37 60.78 44.37 27.63 76.04 82.61 70.93 89.40 82.93 86.77 66.58
TSV Merging 43.73 61.40 43.54 27.94 76.44 83.15 71.65 88.53 82.25 86.41 66.50
Iso-C 43.99 61.34 40.91 22.37 76.96 83.33 71.55 87.74 82.10 85.27 65.56
WUDI Merging 41.39 60.11 44.20 21.05 74.36 80.78 71.12 87.96 81.50 85.48 64.80
OptMerge (Ours) 43.76 61.29 44.68 27.63 76.24 82.97 71.48 89.56 82.97 86.42 66.70
@@
Table 7: Model merging vs. Data mixing.
Methods Solving Time GPU Memory
InternVL2.5-1B (Ours) 0.22h 2.62GB
InternVL2.5-1B (Mixed) 25.38h 240GB
Qwen2-VL-7B (Ours) 3.78h 21.97GB
Qwen2-VL-7B (Mixed) 24.56h 256GB
Computational requirements. As illustrated in Table˜7, we compare the solving time and GPU memory usage of our approach against mixture training. Our approach optimizes the merged vector over 300 iterations while incurring minimal computational overhead and requiring significantly less GPU memory than data-based training. This efficiency is achieved through layer-by-layer optimization without requiring training data. Our results confirm that the proposed method is computationally efficient and highly scalable on devices with modern GPUs, facilitating the rapid development of new models based on existing ones.
Actual checkpoints from Hugging Face. To evaluate the practicality of model merging in communities, we collect fine-tuned models released by different developers on Hugging Face. Our collection includes a model specialized in math reasoning via multimodal reinforcement learning [^2] https://huggingface.co/lmms-lab/Qwen2-VL-7B-GRPO-8k, a personalized model for the Pokemon domain [^3] https://huggingface.co/hiyouga/Qwen2-VL-7B-Pokemon, a model focused on converting PDF documents into text [^4] https://huggingface.co/allenai/olmOCR-7B-0225-preview, and a model with OCR and VQA capabilities in Vietnamese [^5] https://huggingface.co/erax-ai/EraX-VL-7B-V1.0. As shown in Table˜6, OptMerge achieves performance that surpasses that of the individual models, effectively integrating knowledge from diverse models to construct a more robust system.